# Tension-induced giant actuation in unstructured elastic sheets

### v1

@@ Abstract
Note that the scaling law for the width of the curved region, (5), is identical to the scaling for the wavelength of wrinkling observed when a sheet has its whole width clamped and is then pulled perpendicular to this width Cerda and Mahadevan (2003). We will discuss the similarities and the differences between our results and those of Cerda and Mahadevan (2003) (CM) later in the text.
@@
Geometry gives that the inclination angle satisfies $\alpha\approx\tan\alpha=\pi\delta/\omega$ in which $\omega$ has been determined (at least to a scaling level) by the previous argument and we assume $\alpha\ll 1$ for consistency with our use of the FvK formalism; the final part of the puzzle therefore is to determine the maximum amplitude $\delta$ in (4). We note that the imposed mean strain, $\bar{\epsilon}$, is, from the definition (1), $\bar{\epsilon}=\tfrac{\Delta u_{y}}{L}+L^{-1}\int_{0}^{L}\tfrac{1}{2}\left(\partial\zeta/\partial y\right)^{2}~{}\mathrm{d}y$ and, further, that our numerics shows that, even after onset, $\bar{\epsilon}\propto\Delta u_{y}/L$. As a result, we have that $\delta^{2}/L^{2}\propto\bar{\epsilon}-\bar{\epsilon}_{c}$ and $\delta\propto L(\bar{\epsilon}-\bar{\epsilon}_{c})^{1/2}$, which, when combined with the geometrical relationship $\alpha\approx\pi\delta/\omega$, gives us that
@@
Note that here we have employed an empirical observation, based on our numerics, to understand the scaling of the amplitude $\delta$ with strain. Beyond threshold, the non-linear coupling between strain and vertical displacement hinders analytical progress such that even approximate theories based on the FvK framework, e.g. tension field theory (see e.g. the work of Xin and Davidovitch (2021b) on the CM problem Cerda and Mahadevan (2003)), ultimately require numerical resolution.
Figure 3: Master curve for folding angle, $\alpha$ (), as a function of mean strain, $\bar{\epsilon}$. Here, results are shown for both simulations and experiments with rescalings suggested by the critical strain at buckling, (3), and the scaling for folding angle (6). (The experimentally measured strain $\epsilon_{\mathrm{loc}}$ has been converted to an equivalent mean strain based on simulation results — see Fig. A4.)
@@
To highlight the prediction that $\alpha\to 0$ at the critical strain $\bar{\epsilon}_{c}$, we rescale the $x$-axis in Fig. 3 on the basis of the predicted scaling for $\bar{\epsilon}_{c}$ given in (3); this, in turn, suggests a rescaling for the inclination angle. The results plotted in this way show a reasonable collapse (see Fig. 3); a similar collapse is observed when plotting $\epsilon_{\mathrm{loc}}$ in the experimental results (see Fig. A4 of the SI).
@@
Figure 4: Critical buckling behaviour. a, The folding angle $\alpha$ allows for a clearer determination of the critical strain for wrinkling. The model shown here is the same as in Fig. 2a. b, Buckling occurs from where the linear region of the scaled folding angle intersects with the $x$-axis, i.e., at critical buckling strains $\epsilon_{c}$. The curves shown all have $L/W=2$ and $W_{\mathrm{load}}/W=0.2$ with various thickness ratios, $t/W$, indicated by the color bar. The critical strains for various geometric parameters are linearly related to the ratios $(t/W)^{2}$ as shown in the inset.
@@
The prediction for the inclination angle also gives us a more precise means to determine the threshold strain for buckling: except very close to the buckling transition, we expect $\alpha\propto(\bar{\epsilon}-\bar{\epsilon}_{c})^{3/4}$ and so plot $\alpha^{4/3}$ as a function of $\bar{\epsilon}$ (see Fig. 4a). This plot shows the expected linear behaviour and also allows us to determine the threshold strain, $\bar{\epsilon}_{c}$, as a function of the sheet thickness ratio, $t/W$, and Poisson ratio $\nu$. The resulting plot (Fig. 4b) shows that the prediction of (3) is borne out in our numerical simulations, although points with $L=W$ do not collapse as well as those points with $L/W\gg 1$.
@@
Comparing the critical buckling behaviour in our problem (Fig. 4b) and in models of the CM problem (e.g. Fig.7A in Xin and Davidovitch (2021a)), we note that both have identical scaling behaviour with thickness and width of the sheet (both cases are Euler-like buckling) but find that the threshold value of the dimensionless control parameter $\bar{\epsilon}_{c}$ is smaller than $T_{c}^{\text{(CM)}}/Y$ in the CM problem (see ref. Xin and Davidovitch (2021a)). In particular, for $\nu=0.32$ and $t/W=10^{-4}$, we may write the critical mean strain as
@@
$\frac{T_{c}}{Y}\approx\frac{B}{YW^{2}}\times\begin{cases}50,\quad\qquad W_{\mathrm{load}}\ll W,\\ 9\times 10^{4}\quad W_{\mathrm{load}}=W.\end{cases}$ (7)
@@
Note that with the load applied across the full width, the threshold strain required to trigger buckling is more than $1000$ times larger than in the localized case considered here. To understand this enhanced resistance to buckling in the CM problem, we compare the maximum compressive stress in the full-width setup (as calculated in ref. Xin and Davidovitch (2021a)) with that in our problem (see Appendix D). For the specific case of $L/W=8$ and $\nu=0.32$, we find (data not shown) that
@@
$\max_{y}\left\{\frac{-\sigma_{xx}}{T}\right\}\approx\begin{cases}0.5,\quad\qquad W_{\mathrm{load}}\ll W,\\ 0.005,\qquad W_{\mathrm{load}}=W.\end{cases}$ (8)
@@
The magnitude of the maximum compressive stress is hence $100$ times larger in the localized case considered here than in the full-width case. Recalling that the critical strain was $1000$ times larger for the full-width case, we deduce that much of the difference in critical imposed strain can be attributed to the increased efficiency of focusing the longitudinal tensile strain into a transverse compression when pulling at a localized load than at the full-width. However, there also remains an effect of the boundary conditions on the relevant Euler buckling load — indeed, even a standard buckling analysis gives an increase of a factor of 4 between the critical load for a simply-supported beam (analogous to the freer edges here) and a clamped beam (analogous to the more constrained edges of the CM problem) Landau and Lifshitz (1986).
@@
We also found that the scaling for the width of the curved region (5) is equivalent to the wavelength of wrinkles in CM Cerda and Mahadevan (2003). While we emphasize that, for the narrow loading $W_{\mathrm{load}}\ll\omega$ we studied here, only a single depression is formed (equivalent to half a wavelength), this equivalence suggests that the narrow curved strip is effectively wrinkling around the narrowly loaded part. If we relax the assumption $W_{\mathrm{load}}/\omega\ll 1$ — by imposing longitudinal strains large enough that $\omega\sim W_{\mathrm{load}}$, for example — we observe more wrinkle wavelengths in the curved part of the sheet (data not shown). However, one important difference between the localized-loaded sheet and the fully-clamped sheet is that the latter features fully-developed wrinkle patterns in the small-strain regime.
@@
In conclusion, we have shown that tensile loading of an unstructured elastic sheet along one axis actuates motion in an orthogonal direction that is driven by compression. A particularly striking feature of this transformation from tension to compression is that it allows a very small strain to actuate a large change in angle: what we might call ‘Tension-indUced Giant actuation’, or ‘TUG actuation’. We claim that a similar TUG actuation has been observed, if not directly reported, in structured materials — such as a crumpled sheet of paper, kirigami shells Yang et al. (2021), and groovy plastic sheets Meeussen and van Hecke (2023) — but to our knowledge has not been reported in regular materials. While the scaling laws we have derived for the critical strain and the inclination angle reveal how the geometry of the sheet and the properties of the material should be tuned to actuate a homogeneous sheet, in the same way, they also give new insights into the mechanisms by which metasheets are able to achieve TUG actuation. Considering the examples of a crumpled sheet of paper, the kirigami grippers Yang et al. (2021), and the groovy plastic sheet Meeussen and van Hecke (2023), we note that all involve structures (albeit different in nature) that make stretching in one direction easier while bending in the transverse direction is more difficult. For example, the ridges of a crumpled piece of paper buffer excess length Vella (2019) (reducing the effective $Y$), while also increasing the effective thickness (and hence increasing the effective $B$ Košmrlj and Nelson (2017b)), and similarly for the perpendicular cuts and grooves in kirigami grippers and groovy metasheets, respectively. In an unstructured sheet, the combination of a decreased stretching modulus and an increased bending modulus results in an increased buckling strain — see  (2). However, at a given multiple of the strain threshold, $\tilde{\epsilon}=\bar{\epsilon}/\bar{\epsilon}_{c}$, this increase in $B/(YW^{2})$ becomes helpful in achieving TUG actuation; in particular, we note that (6) may be rewritten:
$\alpha\propto\left(\frac{B}{YW^{2}}\right)^{1/4}\frac{L}{W}\tilde{\epsilon}^{1/4}(\tilde{\epsilon}-1)^{1/2},$ (9)
@@
and so, for a given $\tilde{\epsilon}$, large $B/(YW^{2})$, combined with $L/W\gg 1$, leads to a large actuation angle. A second key ingredient therefore is the geometrical focusing that reduces the critical $\bar{\epsilon}_{c}$ by a factor of $\sim 10^{3}$ compared to a fully-clamped case (7). The efficiency of this conversion from tension to transverse compression means that a given strain leads to a large $\alpha$; in addition, if $Y$ is decreased by the structure, such as crumples or cuts, then a fixed load will lead to a still larger strain and hence even larger actuation. This fundamental insight opens the door for designing structures that can more readily demonstrate TUG actuation with or without structure.
@@ Acknowledgments
This work was partially supported by the UK Engineering and Physical Sciences Research Council via Grant No. EP/Y027949/1 (MS and DV); the Ministry of Education, Singapore, via MOE AcRF Tier 3 Award MOE-MOET32022-0002 (KJH); the Presidential Postdoctoral Fellowship from Nanyang Technological University, Singapore, and the start-up funding from the University of Birmingham, UK (ML). We thank Ms Mariona Caus for assisting in taking the pictures in Fig. 1c.
Methods
@@
Test specimens were prepared from commercial PVC films with thickness $t\in\{40,100,150\}\mathrm{~{}\mu m}$. Rectangular sheets were laser cut with width $W=40\mathrm{~{}mm}$, length $L\in\{40,80,200\}\mathrm{~{}mm}$, and two tabs with width $W_{\mathrm{load}}=10\mathrm{~{}mm}$ and length $L_{\mathrm{load}}=30\mathrm{~{}mm}$ at each end, as shown in Fig. A1b. The fixtures were then attached to the blocks of two linear-stage motors that were controlled manually via the knobs. The specimens were clamped at the tabs in 3D-printed clamp fixtures designed with teeth on the inner surfaces to reduce slipping. Prior to clamping the specimen, strain gauges (Tokyo Measuring Instruments Laboratory, GFLA-6-350-70) (which have a gauge length of $6\mathrm{~{}mm}$, resistance of $350\Omega$, and resolution of $1\mu\epsilon$) were adhered to the center of the specimen using strain gauge adhesives (Tokyo Measuring Instruments Laboratory). Strain data measured were recorded by a data logger (Tokyo Measuring Instruments Laboratory, TDS-303). Folding angles of the sheets were measured from photographs taken from the orthogonal plane (Fig. A1d) with the assistance of cannula needles adhered to the edges of the sheet. The needle lines were detected using MATLAB scripts, where the folding angles were measured as the average values of the left and right sides.
@@
Numerical simulation was conducted using the commercial finite element analysis software ABAQUS/2023 using S4R shell elements. Linear elastic constitutive relation was applied with Young’s modulus $E=3000\mathrm{~{}MPa}$ and Poisson’s ratio $\nu=0.3$ (We also varied $\nu$ from 0.01 to 0.49 for the results in Fig. 4). The boundary conditions were set as in the experiment: the top edge of the upper tab was fully fixed, while displacement loading was applied along $y$-axis to the bottom edge of the lower tab. Two analysis steps were conducted in sequence: first, a small out-of-plane displacement were applied to the midline along $x=0$ with a magnitude of $\sim$ 1% of shell thickness, introducing initial imperfections to the planar shell; second, the static analysis step was performed for the process of tensile stretching and sheet folding.

### v2

@@ Abstract
Note that the scaling law for the width of the curved region, (9), is identical to the scaling for the wavelength of wrinkling observed when a sheet has its whole width clamped and is then pulled perpendicular to this width, as studied by CM Cerda and Mahadevan (2003). We will discuss the similarities and the differences between our results and CM later in the text.
@@
Geometry gives that the inclination angle satisfies $\alpha\approx\tan\alpha=\pi\delta/\omega$ in which $\omega$ has been determined (at least to a scaling level) by the previous argument and we assume $\alpha\ll 1$ for consistency with our use of the FvK formalism; the final part of the puzzle therefore is to determine the maximum amplitude $\delta$ in (7). We note that our numerics show that, even after onset, $\bar{\epsilon}\propto\epsilon_{0}$ (see Fig. S6b of the SI Appendix). As a result, we have that $\delta^{2}/L^{2}\propto\bar{\epsilon}-\bar{\epsilon}_{c}$ and $\delta\propto L(\bar{\epsilon}-\bar{\epsilon}_{c})^{1/2}$, which, when combined with the geometrical relationship $\alpha\approx\pi\delta/\omega$, gives us that
@@
Using (6), this result may also be written
$\alpha\propto\frac{tL^{1/2}}{W^{3/2}\sqrt{1-\nu^{2}}}\left(\frac{\bar{\epsilon}}{\bar{\epsilon}_{c}}\right)^{1/4}\left(\frac{\bar{\epsilon}}{\bar{\epsilon}_{c}}-1\right)^{1/2}.$ (11)
@@
Here, we have employed an empirical observation, based on our numerics, to understand the scaling of the amplitude $\delta$ with strain. Beyond threshold, the non-linear coupling between strain and vertical displacement hinders analytical progress such that even approximate theories based on the FvK framework, e.g. tension field theory (see e.g. the work of Xin and Davidovitch Xin and Davidovitch (2021b) on the CM problem Cerda and Mahadevan (2003)), ultimately require numerical resolution.
Figure 3: Master curve for inclination angle, $\alpha$ (), as a function of mean strain, $\bar{\epsilon}$. Here, results are shown for both simulations and experiments with rescalings suggested by the critical strain at buckling, (6), and the scaling for inclination angle (11). The dashed line shows the expected scaling behavior for $\bar{\epsilon}\gg\bar{\epsilon}_{c}$ while the solid curve shows the result (11) assuming a prefactor of $43$ in (6). (The same legends as in Fig. 2b are adopted for simulation data, i.e. marker shapes denote the aspect ratio, marker fill indicates the Poisson’s ratio, marker line-widths denote $W_{\mathrm{load}}/W$ and color indicates the sheet thickness as per the color bar.) Experimental data points are shown behind the simulation data in black-bordered markers with fill colors corresponding to the sheet thickness and shapes denoting the aspect ratio. (The experimentally measured strain $\epsilon_{\mathrm{loc}}$ has been converted to an effective mean strain $\bar{\epsilon}_{\mathrm{eff}}$ based on simulation results — see the plot of $\epsilon_{\mathrm{loc}}$ vs $\bar{\epsilon}$ in Fig. S5 in the SI Appendix. Moreover, to avoid cluttering the plot, prior to buckling every second data point from each simulation is displayed while in the post-buckling regim every eighth data point is displayed.)
@@
To highlight the trend that $\alpha\to 0$ as $\bar{\epsilon}\searrow\bar{\epsilon}_{c}$, we rescale the $x$-axis in Fig. 3 on the basis of the predicted scaling for $\bar{\epsilon}_{c}$ given in (6); (10), in turn, suggests a rescaling for the inclination angle on the $y$-axis. The results plotted in this way show a reasonable collapse (see Fig. 3). The experimental data points in this figure are plotted using an effective mean strain obtained by the linear relation in Fig. S5a to keep a consistent strain definition with the numerics; a similar collapse is observed when plotting $\epsilon_{\mathrm{loc}}$ in the experimental results (see the collapse of the inclination angle in Fig. S5b in the SI Appendix).
@@
Figure 4: Critical buckling behavior in FEM simulations. a The inclination angle $\alpha$ allows for a clearer determination of the critical strain for buckling in our FEM simulations. Buckling occurs where the linear region of the scaled inclination angle intersects with the $x$-axis, which defines the critical buckling strain $\bar{\epsilon}_{c}$. The curves shown all have $L/W=2$, $W_{\mathrm{load}}/W=1/5$, with various sheet thickness ratios as indicated by the color bar. b The critical strains for various geometric parameters and Poisson’s ratios $\nu$ are proportional to the ratio $B/(YW^{2})\propto(t/W)^{2}/(1-\nu^{2})$; the dashed line shows the relationship (6) with coefficient of proportionality $43$.
Critical strain for buckling
@@
The prediction for the inclination angle also gives us a more precise means to determine the threshold strain for buckling in our FEM simulations: except very close to the buckling transition, we expect $\alpha\propto(\bar{\epsilon}-\bar{\epsilon}_{c})^{3/4}$ and so plot $\alpha^{4/3}$ as a function of $\bar{\epsilon}$ (see Fig. 4a). This plot shows the expected linear behavior and also allows us to determine the threshold strain, $\bar{\epsilon}_{c}$, as a function of the sheet thickness ratio, $t/W$, and Poisson’s ratio $\nu$. The resulting plot (Fig. 4b) shows that the prediction of (6) is borne out in our numerical simulations, although points with $L=W$ do not collapse as well as those points with $L/W\gg 1$. We also note that extrapolating this critical strain to the experimentally relevant values of $t/W$ suggests the typical onset strain in experiments is around $10^{-4}$, which for a sample with $L=10\mathrm{~{}cm}$, corresponds to a displacement $\Delta u_{y}\approx 10\mathrm{~{}\mu m}$. This, in addition to the very small angles close to onset already noted, explains why this threshold behavior is not visible in our experiments.
@@
Comparing the critical buckling behavior in our problem (Fig. 4b) and in models of the CM problem (e.g. Fig. 7A of Xin and Davidovitch (2021a)), we note that both have identical scaling behavior with thickness and width of the sheet. However, the effect described here persists as the Poisson effect vanishes, $\nu\to 0$, in contrast to the CM problem. Moreover, we find that the threshold value of the dimensionless control parameter $\bar{\epsilon}_{c}$ is smaller than that in the CM problem Xin and Davidovitch (2021a); for example, comparing the critical strains with $\nu=0.32$ and $t/W=10^{-4}$ we have
@@
$\frac{T_{0}^{c}}{Y}\approx\frac{t^{2}}{W^{2}}\times\begin{cases}43/(1-\nu^{2}),\quad\qquad W_{\mathrm{load}}\ll W,\\ 9\times 10^{4}\quad W_{\mathrm{load}}=W.\end{cases}$ (12)
@@
(Here, the critical value for the full-width case, $W_{\mathrm{load}}=W$, comes from Fig. 7A of Xin and Davidovitch (2021a), while the critical value for $W_{\mathrm{load}}\ll W$ comes from our numerical results in Fig. 4b.) Note that with the load applied across the full width, the threshold strain required to trigger buckling is more than $1000$ times larger than in the localized case considered here. To understand this apparently enhanced resistance to buckling in the CM problem, we compare the maximum compressive stress in the full-width setup (as calculated in ref. Xin and Davidovitch (2021a)) with that in our problem (see the details of the mathematical model in the SI Appendix). For the specific case of $L/W=4$ and $\nu=0.3$, we find (see Fig. S12 of the SI Appendix) that
@@
$\max_{y}\left\{\frac{-\sigma_{xx}}{T_{0}}\right\}\approx\begin{cases}0.6,\quad\qquad W_{\mathrm{load}}\ll W,\\ 0.006,\qquad W_{\mathrm{load}}=W.\end{cases}$ (13)
@@
The magnitude of the maximum compressive stress is hence $100$ times larger in the localized case considered here than in the full-width case. Recalling that the critical strain was $1000$ times larger for the full-width case, we deduce that much of the difference in critical imposed strain can be attributed to the increased efficiency of focusing the longitudinal tensile strain into a transverse compression when pulling at a localized load than with full-width. However, there also remains an effect of the boundary conditions on the relevant Euler buckling load — indeed, even a standard buckling analysis gives an increase of a factor of 4 between the critical load for a simply-supported beam (analogous to the freer edges here) and a clamped beam (analogous to the more constrained edges of the CM problem) Landau and Lifshitz (1986).
@@
We also found that the scaling for the width of the curved region (9) is equivalent to the wavelength of wrinkles in CM Cerda and Mahadevan (2003). While we emphasize that, for the narrow loading $W_{\mathrm{load}}\ll\omega$ we studied here, only a single depression is formed (equivalent to half a wavelength), this equivalence suggests that the narrow curved strip is effectively wrinkling around the narrowly loaded part. If we relax the assumption $W_{\mathrm{load}}/\omega\ll 1$ — by imposing longitudinal strains large enough that $\omega\sim W_{\mathrm{load}}$, for example — we observe wrinkles in the curved part of the sheet (see Fig. S13 of the SI Appendix). However, the localized-loaded sheet and the fully-clamped sheet have significant differences in that localized TUG folding persists as $\nu\to 0$ and the sheet must be loaded to large strains to observe full-width wrinkling.
Conclusions
@@
We have shown that localized tensile loading of an unstructured elastic sheet along one axis actuates folding in an orthogonal direction driven by inducing compression. A particularly striking feature of this transformation from tension to compression is that it allows a very small and localized strain to actuate a giant folding angle: what we call localized ‘TUG folding’. We claim that a similar mechanism has been observed, if not directly reported, in structured materials—such as a crumpled sheet of paper (Fig. 1d), kirigami shells Yang et al. (2021), and groovy plastic sheets Meeussen and van Hecke (2023)—but to our knowledge has not been reported in homogeneous, isotropic materials. While our scaling laws for the critical strain and inclination angle reveal how the geometry of the sheet and the properties of the material should be tuned to actuate a homogeneous sheet, in the same way, they also give new insights into the mechanisms by which the various metasheets (in the examples mentioned above) are able to achieve localized TUG folding. We note that these all involve structures (albeit different in nature) that make stretching in one direction easier but making bending in the transverse direction more difficult. For example, the ridges of a crumpled piece of paper buffer excess length Vella (2019) (reducing the effective $Y$), while also increasing the effective thickness (and hence increasing the effective $B$ Košmrlj and Nelson (2017b)), and similarly for the perpendicular cuts and grooves in kirigami grippers and groovy metasheets, respectively. In an unstructured sheet, the combination of a decreased stretching modulus and an increased bending modulus results in an increased buckling strain — see  (5). However, at a given multiple of the strain threshold, $\tilde{\epsilon}=\bar{\epsilon}/\bar{\epsilon}_{c}$, this increase in $B/(YW^{2})$ becomes helpful in achieving localized TUG folding; in particular, we note that (10) may be rewritten:
$\alpha\propto\left(\frac{B}{YW^{2}}\right)^{1/2}\left(\frac{L}{W}\right)^{1/2}\tilde{\epsilon}^{1/4}(\tilde{\epsilon}-1)^{1/2},$ (14)
@@
and so, for a given $\tilde{\epsilon}$, large $B/(YW^{2})$, combined with $L/W\gg 1$, leads to a large folding angle  [^1] Note that (14) is similar to the result (11) except that for more general, possibly structured, solids, we do not assume that $B\propto Yt^{2}$.. A second key ingredient therefore is the geometrical focusing that reduces the critical $\bar{\epsilon}_{c}$ by a factor of $\sim 10^{3}$ compared to the fully-clamped case (12). The efficiency of this conversion from tension to transverse compression means that a given strain leads to a large $\alpha$; in addition, if $Y$ is decreased by structure, such as cuts, or by the microscopic ‘buffering’ of area Vella (2019), such as crumples or excess area Chen et al. (2022); Valenzuela et al. (2023), then a fixed load will lead to a still larger strain and hence even larger folding angle. This fundamental insight opens the door for designing structures that can more readily demonstrate localized TUG folding with, or without, structure. It may also lead to new insights into similar phenomenology in systems as diverse as dimpled sheets (which fold upon ‘snapping’ of the dimples along a line Liu et al. (2023)) to tension-induced buckling in active systems Fierling et al. (2022).
@@ Acknowledgments
This work was partially supported by the UK Engineering and Physical Sciences Research Council via Grant No. EP/Y027949/1 (MS and DV); the Ministry of Education, Singapore, via MOE AcRF Tier 3 Award MOE-MOET32022-0002 (KJH); the Presidential Postdoctoral Fellowship from Nanyang Technological University, Singapore, and the start-up funding from the University of Birmingham, UK (ML). We thank Ms Mariona Caus for assisting in taking the pictures in Fig. 1d.
Materials and Methods
@@
Test specimens were prepared from commercial PVC films with thickness $t\in\{40,100,150\}\mathrm{~{}\mu m}$. Rectangular sheets were laser cut with width $W=40\mathrm{~{}mm}$, length $L\in\{40,80,200\}\mathrm{~{}mm}$, and two tabs with width $W_{\mathrm{load}}=10\mathrm{~{}mm}$ and length $L_{\mathrm{load}}=30\mathrm{~{}mm}$ at each end (see Fig. S1b of SI Appendix). The fixtures were then attached to the blocks of two linear-stage motors that were controlled manually. The specimens were clamped at the tabs in 3D-printed clamp fixtures designed with teeth on the inner surfaces to reduce slipping. Prior to clamping the specimen, strain gauges with a gauge length of $6\mathrm{~{}mm}$, resistance of $350~{}\Omega$, and resolution of $1~{}\mu\epsilon$ (Tokyo Measuring Instruments Laboratory, GFLA-6-350-70) were adhered to the centre of the specimen using strain gauge adhesives (Tokyo Measuring Instruments Laboratory). The measured strain data were recorded by a data logger (Tokyo Measuring Instruments Laboratory, TDS-303). Inclination angles of the sheets were measured from photographs taken from the orthogonal plane (see the experimental methods in the SI Appendix for further details) with the assistance of cannula needles adhered to the edges of the sheet. The needle lines were detected using MATLAB scripts; the inclination angles reported are the mean values of the left and right sides.
@@
Numerical simulations were conducted using the commercial finite element analysis software ABAQUS/2023. The solid was represented by S4R shell elements and a linear elastic constitutive relation was applied with Young’s modulus $E=3000\mathrm{~{}MPa}$; most simulations were performed with Poisson’s ratio $\nu=0.3$, though results with $\nu=0.01$ and $\nu=0.49$ are also reported. The boundary conditions were set as in the experiment: the top edge of the upper tab was fully fixed, while displacement loading was applied along the $y$-axis to the bottom edge of the lower tab. Two analysis steps were conducted in sequence: first, a small out-of-plane displacement were applied to the midline along $x=0$ with a magnitude of $\sim$ 1% of sheet thickness, introducing initial imperfections to the sheet; second, the static analysis step was performed for the process of tensile stretching and sheet folding.
Supplementary Information